OpenStudy (elisaneedshelp):

If √9x +5=7, what is the value of x? Pleasee Medal and Fan :)

OpenStudy (elisaneedshelp):

\[\sqrt{9x}+5=7\]

geerky42 (geerky42):

subtract both sides by 5 and we have \(\sqrt{9x}=2\), right?

geerky42 (geerky42):

How to get rid of square root? Do you know?

OpenStudy (elisaneedshelp):

No :/ unless i divide both sides by 3x? Im not sure

geerky42 (geerky42):

square root is inverse of square, right? \(2^2=4, \sqrt{4}=2\) \(3^2=9, \sqrt{9}=3\) \(4^2=16, \sqrt{16}=4\) \(\vdots\)

geerky42 (geerky42):

So squaring it will cancel square root and leave us with \(9x\), right?

geerky42 (geerky42):

since we squared \(\sqrt{9x}\), we must also square 2 So we now have \(9x = 4\) Makes sense so far?

OpenStudy (jackthegreatest):

x = 4/9

OpenStudy (jackthegreatest):

that is the next step and also the answer

OpenStudy (jackthegreatest):

hope this helps :)

OpenStudy (bee_see):